OpenStudy (anonymous):

Find all solutions to the equation. sin^(2)x + sin x = 0

OpenStudy (anonymous):

Sin^2x + sinx=0 sinx(sinx+1)= 0 sinx=0 and sinx=-1 , general solution: x=0+n.360 or x= -90 +n.360 or x=270 + n360 ; where n is an element of integers!